OpenStudy (anonymous):

Write a system that describe the situation: Alex purchased 2 hotdogs for x dollars each and 4 popcorn balls for y dollars each for a total cost of $14.00. Marion purchased 4 hotdogs for x dollars each and 7 popcorn balls for y dollars each for a total $26.00.

OpenStudy (anonymous):

2x + 4y = 14 4x + 7y = 26

OpenStudy (anonymous):

Do u need to solve the equation for x & y

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

2x + 4y = 14 -2(2x + 4y = 14) -4x+-8y=-28 4x + 7y = 26 4x + 7y = 26 -y=-2 y=2 2x + 4(2)=14 2x - 8 = 14 2x = 6 x = 3 Hot dogs cost $3 and popcorn balls cost $2 (3,2)

OpenStudy (texaschic101):

nicely done DSS :)

OpenStudy (anonymous):

thnx